# Heterogeneous RBCs via deep multi-agent reinforcement learning

### v1

@@ 3. Experimental setup
In our numerical experiments, we first showcase the capabilities of our framework to reproduce the traditional representative agent RBC model and the mean-field Krusell-Smith (KS) model (Sec. 4.1 and 4.2), and we then show the capabilities of the framework in the general form (Sec. 4.3). In the rest of the paper, we refer to these three experimental settings as ‘RBC’, ‘KS’ and ‘General’.
@@
Parameters. Table 1 reports the main parameters used in our numerical experiments, along with their names and symbols. The values reported are grouped depending on the type of experiment performed, either ‘RBC’, ‘KS’, or ‘General’. If a single value is reported, then the same value has been used for all experiments. We note here that most of the parameters used are standard for macroeconomic models of these types Romer (2018).
@@
Figure 3. MARL-BC can reproduce representative‐agent results with a single agent. The left panel shows how the RL household recovers the optimal choices of consumption fraction $\hat{c}$ and labour supply $\ell$ as given in Eq. (13) for a textbook, analytically solvable, RBC model with full capital depreciation ($\delta=1)$. The centre panel shows a similar convergence, but for a more complex model with a partial depreciation ($\delta=0.025$). This model cannot be analytically solved, and the optimal choices are obtained here using numerical tools common in economics. Finally, the right panel shows that MARL-BC is also capable of reproducing the dynamical behaviour of the standard RBC models as captured by an impulse response function. The plot shows the impulse response for consumption following a productivity shock of one standard deviation, with the dashed orange line indicating the standard RBC result and the full blue line indicating the equivalent MARL-BC behaviour.
@@
Training and testing. We implemented MARL-BC using well-known open source libraries in Python. Specifically, we developed our MARL environment following the interface offered by PettingZoo Terry et al. (2021), and use SuperSuit Terry et al. (2020) to use RL algorithms from Stable-Baselines3 Raffin et al. (2021) in an independent MARL fashion. We consider the following RL algorithms for the household agents: Deep Deterministic Policy Gradient (DDPG) Lillicrap et al. (2016a), Twin Delayed Deep Deterministic policy gradient (TD3) Fujimoto et al. (2018a), Soft Actor Critic (SAC) Haarnoja et al. (2018a), Proximal Policy Optimization (PPO) Schulman et al. (2017).
@@
To facilitate quick convergence and scalability, we adopt the standard MARL paradigm of parameter sharing Albrecht et al. (2024). In practice, this means the neural networks representing the different agents in the economy share the same parameters or, equivalently, that a single neural network represents all the agents. Importantly, this paradigm does not force an identical behaviour for all agents, since the different agents’ characteristics (such as the individual productivities of labour and capital) appear as observations for the neural network, which can thus encode completely different behaviours. Parameter sharing significantly reduces the computational burden, improves sample efficiency and scalability and can help to stabilise training. Furthermore, it can also be thought of as a stylised form of ‘social learning’ Whalen et al. (2018); Ndousse et al. (2021) since information discovered by an agent is immediately made available to similar agents via shared parameters.
@@
We train all algorithms for $10^{6}$ steps in the single-agent environment and note that $10^{5}$ steps are typically sufficient for convergence. Hence, for the multi-agent runs we perform $10^{5}$ per-agent updates, corresponding to $n\cdot 10^{5}$ total steps. The code for the MARL-BC environment used to perform our experiments is available in open source at https://github.com/fedegabriele/MARL-BC.
@@ 4. Results
Figure 4. MARL-BC can reproduce mean‐field Krusell–Smith results with a large number of identical agents. The left panels are scatter plots of the aggregate capital at two consecutive times $K_{t}$ and $K_{t+1}$, at four training snapshots, with progressively more training steps (namely $2^{7}$, $2^{11}$, $2^{19}$ and $2^{21}$, which approximately amount to $10^{2}$, $2\cdot 10^{3}$, $5\cdot 10^{4}$, and $2\cdot 10^{6}$). A characteristic result of the KS model is the linear time evolution of aggregate capital (as shown in the first figure of the original paper Krusell and Smith (1998)), which is here recovered with a sufficient training steps. The centre panels depict the wealth distributions of the agents before and after training has converged with a sufficiently large number of training steps. While the Gini index for the untrained model is 0.08, it moves to 0.18 for the trained model distribution, a number which is in line with the original work. Finally, the right panels illustrate the curves of the marginal propensities to consume learned by the MARL-BC agents. Analogously to the KS agent policy, the consumption policy learned by the MARL-BC agents is almost completely flat when wealth is high, while it increases sharply when wealth is low.
@@ 4.1. Representative agent RBC limit
In this way, individual and aggregate quantities coincide: $K_{t}=k_{t}$ and $L_{t}=\ell_{t}$. Furthermore, as for all standard RBCs, we choose the technology to move according to the exponential of an AR(1) process, i.e.
@@
Reproducing textbook RBCs ($\boldsymbol{\delta=1}$). Here, we reproduce the results of the simplest type of RBC model as covered in most macroeconomic textbooks, one with full depreciation ($\delta=1$). In this setting, capital cannot be transferred from one step to the following, and closed-form solutions can be obtained Brock and Mirman (1972). The optimal policy $(\hat{c}^{\star}_{t},\ell^{\star}_{t})$ reads
@@
Reproducing typical RBCs ($\boldsymbol{\delta=0.025}$). More general RBCs set up a partial depreciation, often of $\delta=0.025$ as done in our experiments. Relaxing the full-depreciation assumption makes the model impossible to solve analytically. In these cases, economic modellers typically resort to approximate solutions relying on first-order conditions and linearisation as implemented in widely spread software such as “Dynare”Villemot et al. (2011). The centre panel of Figure 3 shows that, after approximately $10^{4}-10^{5}$ training steps, the RL household of our framework learns optimal consumption and labour choices that are coherent with those computed with the Dynare software.
@@
Figure 5. MARL-BC allows for modelling rich ex-ante heterogeneity and for the emergence of heterogeneous behaviour. The left panel reports the Lorenz curves, and the corresponding Gini indices, of the wealth distributions computed using the MARL-BC framework, either with the standard KS setting (blue dots) or with the extended KS with heterogeneous capital productivities (orange circles and green triangles). The graph illustrates how MARL-BC allows for the modelling of a much wider range of wealth inequalities. The centre panel reports the marginal propensities to consume of the three classes of agents (of low, medium and high returns) for the KS model with the Gini index equal to 0.61. The graph illustrates how MARL-BC gives rise to very different choices for the three household groups. The right panel illustrates the kind of wealth distribution that one can obtain by setting up an RBC model with 9 agents, each with a distinct combination of labour and capital productivity. Higher productivities also imply higher wages or capital returns that, in turn, give rise to higher capital accumulation.
4.2. Mean-field Krusell-Smith limit
The celebrated ‘Krusell-Smith’ (KS) model Krusell and Smith (1998) is a mean field model with ex-ante identical households that face discrete shocks of individual unemployment and aggregate technology. As illustrated in the bottom right panel of Figure 1, our framework can recover the KS model with a large number of identical agents and appropriately defined shocks. In other words, our approach can give rise to mean field results in the limit
@@ 4.2. Mean-field Krusell-Smith limit
In practice, in our experiments we use $n=20$ agents as we expect this to be sufficient to reproduce trends and distributions of the mean field limit. Furthermore, to recover the precise KS model behaviour, we set up the aggregate and individual shocks in close agreement with the original model. Specifically, we let the aggregate technology factor $A_{t}$ follow a two‐state Markov chain $A_{t}\in[A_{t}^{g},A_{t}^{b}]$. In the “good” ($g$) or ”bad” ($b$) state, technology is respectively higher $A_{t}^{g}=1.02$ or lower $A_{t}^{b}=0.98$. Furthermore, individual households also follow a two-state Markov chain, being either employed and supplying $\ell^{i}_{t}=\bar{\ell}=1.11$ units of labour, or being unemployed and supplying no labour ($\ell^{i}_{t}=0$). The transition probabilities between aggregate and individual states are taken from Krusell and Smith (1998). They are calibrated to make the expected duration of each regime eight periods, and in such a way that the unemployment rate is around $4\%$ in good states and $10\%$ in bad states. Since labour is fixed, we let the RL household choose only the consumption fraction $\hat{c}^{i}_{t}$ and set $b=0$ in the reward function. Finally, in agreement with Krusell and Smith (1998), we set the observation space to $\mathbf{x}^{i}_{t}=(k^{i}_{t},\ell^{i}_{t-1},K_{t},A_{t})$.
@@
Reproducing the classic KS law of motion. The left panel of Figure 4 illustrates the convergence properties of our MARL-BC framework relative to the classic ‘law of motion’ for aggregate capital postulated in the original paper Krusell and Smith (1998). The four small scatter plots display the pairs $(K_{t},K_{t+1})$ after approximately $10^{2}$, $2\cdot 10^{3}$, $5\cdot 10^{4}$, and $2\cdot 10^{6}$ SAC updates. In the KS procedure, a linear map is postulated whose coefficients are estimated until they converge in a self-consistent procedure. By contrast, using MARL-BC a perfectly linear relationship ($R^{2}>0.99$) between consecutive aggregate capitals emerges endogenously without any a priori assumption.
@@
Reproducing KS distributional properties. The two histograms in the centre panel Figure 4 report the stationary distribution of individual capital before and after convergence. Before convergence, the scale of the capital distribution is very small, ranging only from 0 to 1.2. This is a sign that the agents have not yet learned consumption policies capable of accumulating capital. Not surprisingly, the Gini index is very low. After convergence, the Gini index increases to $0.18$, a value relatively close to the 0.25 value from the original KS calculation.
@@
Reproducing KS marginal propensity to consume. The two graphs in the right panel Figure 4 report the agents’ marginal propensity to consume as a function of their wealth, before and after training (top and bottom graphs). The graphs show that the learning process converges to curves that are flat for large values of wealth ($a>20$) and increase rapidly for low values of wealth (especially after $a<5$). Interestingly, there is very little difference in the consumption fractions if the agents are employed (orange dots) or unemployed (blue dots). These shapes for the curves of the marginal propensity to consume are another key result of the original KS paper that the MARL-BC can easily recover.
@@ 4.3. Greater heterogeneity with MARL-BC
We here illustrate how MARL-BC can seamlessly extend classic models by leveraging its flexibility to represent a greater level of agent heterogeneity. Specifically, we extend the KS model with heterogeneous capital productivities and then extend the RBC model with heterogeneity both in capital and labour productivities.
KS with heterogeneous capital returns. We extend the KS by dividing the $n=20$ agents into three groups, respectively with low, middle and high capital productivity. Specifically, we assign 14 agents (or 70% of the total) to the middle productivity group, and 6 agents (or 30% of the total) equally between the low and high productivity groups. Although this assignment is largely arbitrary, the described proportions were selected to resemble those of a normal distribution, where roughly 70% of the data falls within one standard deviation of the mean and 30% lies in the two tails. We perform two experiments, with either mild or marked differences in capital productivity among the three groups. To be more precise, in the ‘mild’ experiment we let the three productivities $\kappa^{i}$ be $\{0.8,1.0,1.2\}$, while in the ‘marked’ experiment we let them be $\{0.0,1.0,1.2\}$. As capital productivities have an immediate effect on capital returns via Eq. (3), this KS extension can neatly represent the heterogeneous returns on capital of real economies Xavier (2021).
@@
RBC with heterogeneous returns and wages. To further illustrate the flexibility and scope of MARL-BC, we build a heterogeneous version of the standard RBC described in Sec. 4.1 with nine agents, each with capital and labour productivity taken from a $3\times 3$ grid containing values in the set $\{0.98,1,1.02\}$. So that, the agent with the lowest productivity has $\kappa^{i}=\lambda^{i}=0.8$, the one with the highest productivity has $\kappa^{i}=\lambda^{i}=1.2$, and the other agents have all combinations in the middle. In turn, these different productivities give rise to heterogeneity in both capital returns and wages, which influence consumption and labour decisions as well as wealth levels. In the right panel of Figure 5, we illustrate the results of this setup by showing how the nine agents stabilise with overlapping yet distinct levels of wealth after a sufficient number of training steps. Although we do not show this in the figure, the different productivities also give rise to significantly different choices for consumption fraction and labour supply.
Scalability to hundreds of agents. To test the scalability of MARL-BC to simulations with a large number of heterogeneous agents, we expand on the experiment discussed in the previous paragraph and perform training runs for progressively denser, equispaced, grids of capital and labour productivities ($\kappa^{i}$ and $\lambda^{i}$), from $3\times 3=9$ agents to $23\times 23=529$ agents. The results of these runs are shown in Figure 6. The top panel in the figure, reports mean and two standard deviations of the best episodic return attained during training runs of $10^{5}$ per-agent updates. The bottom panel reports, the corresponding wall-clock time taken for training on a single CPU. We find that the performance of SAC remains stably high across model sizes, from tens to hundreds of agents. On the contrary, we find that PPO and DDPG strongly underperform in regimes of low and high number of agents respectively, as compared to their competitors. Finally, we find that the computational cost rises steadily with the number of agents, and yet remains within practical limits for SAC (it takes roughly two hour to train the largest model for roughly $50\cdot 10^{6}$ steps). These results indicate that MARL-BC can be successfully trained with hundreds of heterogeneous households, even using very modest hardware.

### v2

@@ 3. Experimental setup
In our numerical experiments, we first showcase the capabilities of our framework to reproduce the traditional representative agent RBC model and the mean-field Krusell–Smith (KS) model (Sec. 4.1 and 4.2), and we then show the capabilities of the framework in the general form (Sec. 4.3). In the rest of the paper, we refer to these three experimental settings as ‘RBC’, ‘KS’ and ‘General’.
@@
Parameters. Table 1 reports the main parameters used in our numerical experiments, along with their names and symbols. The values reported are grouped depending on the type of experiment performed, either ‘RBC’, ‘KS’, or ‘General’. If a single value is reported, then the same value has been used for all experiments. We note here that most of the parameters used are standard for macroeconomic models of these types romer2018advanced.
@@
Figure 3. MARL-BC can reproduce representative‐agent results with a single agent. The left panel shows how the RL household recovers the optimal consumption fraction $\hat{c}$ and labour supply $\ell$ of Eq. (13), for a textbook, analytically solvable, RBC model with full capital depreciation ($\delta=1)$. The centre panel shows a similar convergence, but for a more complex model with a partial depreciation ($\delta=0.025$). This model cannot be analytically solved, and the optimal choices are obtained here using numerical tools common in economics. Finally, the right panel shows that MARL-BC is also capable of reproducing the dynamical behaviour of the standard RBC models as captured by an impulse response function. The plot shows the impulse response for consumption following a productivity shock of one standard deviation, with the dashed orange line indicating the standard RBC result and the full blue line indicating the equivalent MARL-BC behaviour.
@@
Training and testing. We implemented MARL-BC using well-known open-source libraries in Python. Specifically, we use the MARL interface of PettingZoo terry2021pettingzoo, and use SuperSuit terry2020supersuit to leverage RL algorithms from Stable-Baselines3 raffin2021stable in an independent MARL fashion. We consider the following RL algorithms: Deep Deterministic Policy Gradient (DDPG) LillicrapHPHETS15, Twin Delayed Deep Deterministic Policy Gradient (TD3) fujimoto2018addressing, Soft Actor Critic (SAC) haarnoja2018soft, Proximal Policy Optimization (PPO) schulman2017proximal. We use default hyperparameters as detailed in (gabriele2025heterogeneous_ARXIV, Appendix).
@@
To facilitate quick convergence and scalability, we adopt the standard MARL paradigm of parameter sharing albrecht2024multi. In practice, this means a single neural network represents all the agents. Importantly, this paradigm does not force an identical behaviour for all agents, since the different agents’ characteristics (such as the individual productivities of labour and capital) appear as observations for the neural network, which can thus encode completely different behaviours. Parameter sharing significantly reduces the computational burden, improves sample efficiency and scalability, and can help to stabilise training. Furthermore, it can also be thought of as a stylised form of ‘social learning’ whalen2018sensitivity; ndousse2021emergent since information discovered by an agent is immediately made available to similar agents via shared parameters.
@@
We train all algorithms for $10^{6}$ steps in the single-agent environment. For the multi-agent runs, we perform $10^{5}$ per-agent updates, corresponding to $n\cdot 10^{5}$ total steps, sometimes more if needed for convergence (as detailed in (gabriele2025heterogeneous_ARXIV, Appendix)). Unless otherwise stated, we report mean and standard deviation over five different training runs. The code for the MARL-BC environment used to perform our experiments is available in open source at https://github.com/fedegabriele/MARL-BC.
@@ 4. Results
Figure 4. MARL-BC can reproduce mean‐field Krusell–Smith results with a large number of identical agents. The left panels are scatter plots of the aggregate capital at two consecutive times $K_{t}$ and $K_{t+1}$, at four training snapshots, with progressively more training steps (namely $2^{7}\approx 10^{2}$, $2^{11}\approx 2\cdot 10^{3}$, $2^{19}\approx 5\cdot 10^{4}$ and $2^{21}\approx 2\cdot 10^{6}$). The characteristic linear time evolution of aggregate capital (cf. the first figure of the original paper KrusellSmith1998) is here recovered with a sufficient number of training steps. The centre panels depict the wealth distributions of the agents before and after training. While the Gini index for the untrained model is 0.08, it moves to 0.18 for the trained model distribution, a number which is in line with the original work. Finally, the right panels illustrate the curves of the marginal propensities to consume learned by the MARL-BC agents. Analogously to the KS agent policy, the consumption policy learned by the MARL-BC agents is almost completely flat when wealth is high, while it increases sharply when wealth is low.
@@ 4.1. Representative agent RBC limit
so that individual and aggregate quantities coincide: $K_{t}=k_{t}$ and $L_{t}=\ell_{t}$. Furthermore, as for all standard RBCs, we choose the technology to move according to the exponential of an AR(1) process
@@
Reproducing textbook RBCs ($\boldsymbol{\delta=1}$). Here, we reproduce the results of the simplest type of RBC model as covered in most macroeconomic textbooks, one with full depreciation ($\delta=1$). In this setting, capital cannot be transferred to the next step and closed-form solutions can be obtained brock1972optimal. The optimal policy $(\hat{c}^{\star}_{t},\ell^{\star}_{t})$ reads
@@
Reproducing typical RBCs ($\boldsymbol{\delta=0.025}$). More general RBCs set up a partial depreciation, often of $\delta=0.025$ as done in our experiments. Relaxing the full-depreciation assumption makes the model impossible to solve analytically. In these cases, economic modellers typically resort to approximate solutions relying on first-order conditions and linearisation as implemented in widespread software such as ‘Dynare’ villemot2011solving. The centre panel of Figure 3 shows that, after approximately $10^{4}-10^{5}$ training steps, the RL household of our framework learns optimal consumption and labour choices that are coherent with those computed with the Dynare software.
@@
Figure 5. MARL-BC allows for modelling rich ex-ante heterogeneity and for the emergence of heterogeneous behaviour. The left panel reports the Lorenz curves, and the corresponding Gini indices, of the wealth distributions computed using MARL-BC, either in the standard KS setting (blue dots) or in the extended KS with heterogeneous capital productivities (orange circles and green triangles). The graph illustrates how MARL-BC allows one to model a wider range of wealth inequalities. The centre panel reports the marginal propensities to consume of the three classes of agents (of low, medium and high returns) for the KS model with Gini index equal to 0.61. The graph illustrates how the three household groups learn markedly different policies. The right panel illustrates the kind of wealth distribution that one can obtain from an RBC model with 9 agents with different combinations of labour and capital productivity. The legend shows only 3 colours, but the pattern is clear: higher productivities imply higher wages or capital returns that, in turn, give rise to higher capital accumulation.
4.2. Mean-field Krusell–Smith limit
The celebrated ‘Krusell–Smith’ (KS) model KrusellSmith1998 is a mean-field model with ex-ante identical households that face discrete shocks of individual unemployment and aggregate technology. As illustrated in the bottom right panel of Figure 1, our framework can recover the mean-field KS limit with a large number of identical agents
@@ 4.2. Mean-field Krusell–Smith limit
In practice, in our experiments we use $n=20$ agents as we expect this to be sufficient to reproduce trends and distributions of the mean field limit. Furthermore, to recover the precise KS model behaviour, we set up the aggregate and individual shocks in close agreement with the original model. Specifically, we let the aggregate technology factor $A_{t}$ follow a two‐state Markov chain $A_{t}\in[A_{t}^{g},A_{t}^{b}]$. In the ‘good’ ($g$) or ‘bad’ ($b$) state, technology is respectively higher $A_{t}^{g}=1.02$ or lower $A_{t}^{b}=0.98$. Furthermore, individual households also follow a two-state Markov chain, being either employed and supplying $\ell^{i}_{t}=\bar{\ell}=1.11$ units of labour, or being unemployed and supplying no labour ($\ell^{i}_{t}=0$). The transition probabilities between aggregate and individual states are taken from KrusellSmith1998. They are calibrated to US data to make the expected duration of each regime eight periods, and in such a way that the unemployment rate is around $4\%$ in good states and $10\%$ in bad states. Since labour is fixed, we let the RL household choose only the consumption fraction $\hat{c}^{i}_{t}$ and set $b=0$ in the reward function. Finally, in agreement with KrusellSmith1998, we set the observation space to $\mathbf{x}^{i}_{t}=(k^{i}_{t},\ell^{i}_{t},K_{t},A_{t})$. The results of these experiments are given in Figure 4 through a single illustrative training run, and discussed below. Statistics over different seeds are available in (gabriele2025heterogeneous_ARXIV, Appendix).
@@
Reproducing the classic KS law of motion. The left panel of Figure 4 illustrates the convergence properties of our MARL-BC framework relative to the classic ‘law of motion’ for aggregate capital postulated in the original paper KrusellSmith1998. The four small scatter plots display the pairs $(K_{t},K_{t+1})$ after approximately $10^{2}$, $2\cdot 10^{3}$, $5\cdot 10^{4}$, and $2\cdot 10^{6}$ SAC updates. In the KS procedure, a linear map is postulated whose coefficients are estimated until they converge in a self-consistent procedure. By contrast, using MARL-BC a perfectly linear relationship ($R^{2}>0.99$) between consecutive aggregate capitals emerges endogenously without any a priori assumption.
@@
Reproducing KS distributional properties. The two histograms in the centre panel of Figure 4 report the stationary distribution of individual capital before and after convergence. Before convergence, the scale of the capital distribution is very small, ranging only from 0 to 1.2. This is a sign that the agents have not yet learned consumption policies capable of accumulating capital. Not surprisingly, the Gini index is very low. After convergence, the Gini index increases to $0.18$, a value relatively close to the 0.25 value from the original KS calculation.
@@
Reproducing KS marginal propensity to consume. The two graphs in the right panel of Figure 4 report the agents’ marginal propensity to consume as a function of their wealth, before and after training (top and bottom graphs). The graphs show that the learning process converges to curves that are flat for large values of wealth ($a>20$) and increase rapidly for low values of wealth (especially after $a<5$). Interestingly, there is very little difference in the consumption fractions if the agents are employed (orange dots) or unemployed (blue dots). These shapes for the curves of the marginal propensity to consume are another key result of the original KS paper that the MARL-BC can easily recover.
@@ 4.3. Greater heterogeneity with MARL-BC
We illustrate here how MARL-BC can seamlessly extend classic models by leveraging its flexibility to represent a greater level of agent heterogeneity. Specifically, we extend the KS model with heterogeneous capital productivities and then extend the RBC model with heterogeneity in both capital and labour productivities.
KS with heterogeneous capital returns. We extend the KS model by dividing the $n=20$ agents into three groups, respectively with low, middle and high capital productivity. Specifically, we assign 14 agents (or 70% of the total) to the middle productivity group, and 6 agents (or 30% of the total) equally between the low and high productivity groups. Although this assignment is largely arbitrary, the described proportions were selected to resemble those of a normal distribution, where roughly 70% of the data falls within one standard deviation of the mean and 30% lies in the two tails. We perform two experiments, with either mild or marked differences in capital productivity among the three groups. To be more precise, in the ‘mild’ experiment we let the three productivities $\kappa^{i}$ be $\{0.8,1.0,1.2\}$, while in the ‘marked’ experiment we let them be $\{0.0,1.0,1.2\}$. As capital productivities have an immediate effect on capital returns via Eq. (3), this KS extension can neatly represent the heterogeneous returns on capital of real economies xavier2021wealth.
@@
RBC with heterogeneous returns and wages. To further illustrate the flexibility and scope of MARL-BC, we build a heterogeneous version of the standard RBC described in Sec. 4.1 with nine agents, each with capital and labour productivity taken from a $3\times 3$ grid containing values in the set $\{0.98,1,1.02\}$. Thus, the agent with the lowest productivity has $\kappa^{i}=\lambda^{i}=0.98$, the one with the highest productivity has $\kappa^{i}=\lambda^{i}=1.02$, and the other agents have all combinations in the middle. In turn, these different productivities give rise to heterogeneity in both capital returns and wages, which influence consumption and labour decisions as well as wealth levels. In the right panel of Figure 5, we illustrate the results of this setup by showing how the nine agents stabilise with overlapping yet distinct levels of wealth after a sufficient number of training steps. Although we do not show this in the figure, the different productivities also give rise to significantly different choices for consumption fraction and labour supply.
Scalability to hundreds of agents. To test the scalability of MARL-BC to simulations with a large number of heterogeneous agents, we expand on the experiment discussed in the previous paragraph and perform training runs for progressively denser, equispaced, grids of capital and labour productivities ($\kappa^{i}$ and $\lambda^{i}$), from $3\times 3=9$ agents to $23\times 23=529$ agents. The results of these runs are shown in Figure 6. The top panel in the figure reports the mean and two standard deviations of the best episodic return attained during training runs of $10^{5}$ per-agent updates. The bottom panel reports the corresponding wall-clock time taken for training on a single CPU. We find that the performance of SAC remains stably high across all model sizes. On the contrary, we find that PPO and DDPG strongly underperform respectively in regimes of low and high numbers of agents. Finally, we find that the computational cost rises steadily with the number of agents, and yet remains within practical limits for SAC (it takes roughly two hours to train the largest model for roughly $50\cdot 10^{6}$ steps). These results indicate that MARL-BC can be successfully trained with hundreds of households, even using very modest hardware. Further details on the scalability of the framework are available in (gabriele2025heterogeneous_ARXIV, Appendix).